# Generating Synthetic Data via Augmentations for Improved Facial Resemblance in DreamBooth and InstantID

### v1

@@ 4.2.2 Landmarks Image
For practical applications, users frequently struggle to understand how face positioning in the landmark image transfers to the generated output. This communication challenge often results in user dissatisfaction with generated images, despite the issue stemming from suboptimal conditioning input.
To address this limitation, we propose two solutions: 2-shot generation and face replacement.
In 2-shot generation, we collect subject reference images $(s_{1},\dots,s_{n})$ and a separate image representing the desired pose and composition $s_{kpts}$. These are used as reference images and the keypoints image, respectively. While the resulting output $s_{out}$ is generally satisfactory, using facial landmarks from one person to generate another reduces facial similarity due to structural differences in the five keypoints (eyes, nose, mouth). We hypothesize this stems from imbalanced conditioning weights. Performance improves when replacing $s_{kpts}$ with a previously generated image of the subject, yielding better facial similarity while maintaining compositional control.
In face replacement, users interact with a simple tool to manipulate (move/rotate/resize) their cropped face on a canvas matching the diffusion model’s output dimensions. This approach eliminates the similarity issues caused by using another person’s facial landmarks. However, the method performs poorly when none of the reference images show the subject facing the camera (deviations $>30$ degrees). User satisfaction was higher with this approach compared to 2-shot generation, which we attribute to increased interactivity and faster generation times.
@@ 5 Survey
We conducted the survey to evaluate the viability of AI-generated portraits for professional use and to compare the performance of DreamBooth and InstantID in generating realistic headshots. 97 white-collar workers from diverse professional backgrounds participated in the online survey. Numerical data can be found in Suppl. 11 and questionary can be found in Suppl. 12.
Overall Performance of Generated Portraits
Portraits generated by DreamBooth and InstantID performed similarly across multiple aspects, including overall quality, facial detail clarity, identity preservation, perceived level of editing, and background quality. Using high-quality subject datasets led to slightly better results in most categories, except for ”Editing,” where participants indicated familiarity and acceptance of traditional Photoshop-enhanced portraits.
Method Preferences
A slightly higher percentage of participants ( 4%) preferred the standardized portraits from InstantID over the more flexible outputs of DreamBooth. InstantID was often perceived as more professional, likely due to its consistent ”Photoshopped look,” which resonated with a broader audience. Open-ended responses highlighted diverse preferences, with participants emphasizing factors such as lighting, pose, angle, expression, detail, color, and background.
Facial Similarity
DreamBooth demonstrated superior facial similarity between real images individuals and their generated portraits compared to InstantID. More participants identified InstantID images as depicting a different person than the reference. DreamBooth consistently maintained a higher level of facial similarity across both high- and low-quality subject datasets.
Noticing AI Generations
Most white-collar workers struggled to identify AI-generated headshots when not explicitly prompted, often focusing on well-known but absent flaws commonly associated with AI generation. Among a subset of participants $(n=77)$ who regularly notice AI-generated images in daily life, the generated portraits blended well with conventional studio photographs. However, participants who actively use AI for image creation $(n=29)$ demonstrated better identification skills. This group was more likely to recognize DreamBooth images as AI-generated, possibly due to DreamBooth’s popularity, while InstantID generations, being more niche, had a near $/50$ 50 chance of being identified as AI.
@@ 6 Discussion
Our experiments offer insights into augmentation strategies for improving facial resemblance in personalized text-to-image generation using DreamBooth and InstantID. While classical augmentations are common in deep learning, applying them to few-shot personalization can yield undesirable results. Geometric transformations like flipping and rotation disrupted learning due to face asymmetry and artifacts. Color jittering caused erratic generations by associating color shifts with DreamBooth’s rare token. Background augmentations with U2-Net introduced segmentation imperfections, especially around hair, which the model learned. Replacing backgrounds with patterns or studio backdrops also degraded image quality. However, auto color adjustment with Adobe Lightroom improved color grading.
@@
FaceDistance provided a quantitative measure of facial similarity but became less useful for hyperparameter tuning once a certain fidelity level was reached. A user survey among white-collar workers showed that both DreamBooth and InstantID performed similarly in quality, clarity, identity preservation, editing, and background. A slight preference emerged for the ”Photoshopped look” of InstantID portraits. While DreamBooth achieved better facial similarity, many participants struggled to distinguish AI-generated images from real ones, particularly those unfamiliar with AI tools. Users actively engaged in AI image creation were more likely to identify DreamBooth images as synthetic, possibly due to its higher popularity.
InstantID’s effectiveness depends on reference image quality and diversity. Using multiple references (around four) improved similarity by enriching information for the Face Encoder. Facial landmarks strongly influenced pose and composition, sometimes overriding text prompts. We explored 2-shot generation and interactive face replacement to enhance control, with the latter showing higher user satisfaction. Rotational and shape-altering augmentations were ineffective, and background modifications reduced similarity. Traditional upscaling worked well for low-resolution images, whereas neural network-based upscaling introduced artifacts.
@@ 8 Conclusion
For practical use, employing multiple reference images enhances facial information capture. Improving control over pose and composition through landmarks is crucial, with interactive face replacement showing promise.
@@ 9 FaceDistance of InstantID generations
Our experiments with InstantID reveal that using a single reference image produces high variance in output quality with inconsistent facial similarity. (Figure 5) When using multiple reference images, results significantly improve facial representation consistency (Figure 6). Analysis of the Vacation-Anna dataset demonstrates that while the visual improvement between using 1 and 2 reference images is substantial, the differences between using 2, 4, or 8 images are subtle and difficult to distinguish without direct side-by-side comparison. The FaceDistance metric confirms this, showing that as more reference images are used, the distribution of distances shifts leftward, indicating better overall similarity.
Figure 5: Distribution of facial similarity metrics for 8 different input subjects using single reference images, with the y-axis sorted by FaceDistance (Img 3 showing lowest inter-image distance, Img 5 showing highest).
@@
10 Motivation to Seek Alternatives to InstantID
@@ 11 Comparative Survey Analysis of Portrait Generation Methods
Our analysis reveals that generated portraits perform similarly across individual aspects, likely due to the SD model employed (Figure 9). Good subject datasets for Dreambooth yield more natural results, though user preferences trend toward Photoshop-edited appearances. Further investigation shows that ”good” datasets correlate with higher employee agreement on preferred methods, with approximately 4% more participants favoring standardized InstantID portraits for their professional, photoshopped aesthetic (Figure 10). When assessing facial similarity between real and generated images, Dreambooth consistently outperformed InstantID regardless of dataset quality, with fewer participants identifying Dreambooth generations as different individuals (Figure 11).
White-collar workers demonstrated limited ability to distinguish AI-generated images, often searching for well-known AI flaws absent from our high-quality generations. Among the subset of participants (n = 77) who reported noticing AI-generated images in daily life, our generations integrated seamlessly with non-generated studio images, with only the ”Real (Daniel)” image—featuring strong contrast and sharp features—dividing opinion nearly equally (Figure 12).
@@
Figure 10: Percentage of Participant Choices.
Figure 11: Percentage of Similarity Ratings per Method.
Figure 12: Percentage of No/Yes Answers per Picture. $(n=77)$
12 Survey Questionare
Here is the full questionnaire for the survey. The answer options are in italic. Titles and other guiding instructions are in bold. We included descriptions to the images.
@@ 1. Title: Survey on Studio Portraits
    Short introduction
@@
    Rate the skill level of the photographer. (This is repeated for all four subjects.) Each subject has four generated images arranged in a grid. A short description of their professional background is included (e.g., marketer, researcher, nurse).
@@
      How would you rate the overall quality? [Really Bad] 1 - 5 [Really Good]
@@
      Are the facial details clear and well-defined? [No] 1 - 5 [Yes]
@@
      . [Completely Different] 1 - 5 [Exactly The Same]
@@
      . [No Editing] 1 - 5 [Heavily Edited]
@@
      Rate the quality of the background in the headshot. [Poor] 1 - 5 [Excelent]
@@
      Would you expect to see these photos in a professional context, such as on LinkedIn or company websites? [Yes/No]
@@
      Is there anything you don’t like about these pictures?      [Text Answer]
@@
    Which photographers were good? If you’ve liked at least 1 image, click the square near it. [Four image grids from above]
@@
    Similarity of Real-Life Pictures and Portraits. In this section, you are asked to compare how similar the person in the everyday pictures on the left is to the person in the portraits on the right.
    How would you rate the similarity of the person? . [Not the same] 1-3 [The Same]
@@
    Detection of the use of AI.  In this chapter, we might have used artificial intelligence (AI) in a way. Don’t go back to previous chapters and change your answers. This is important for our research.
    Please select photos, if any, that you think AI was used. [Four pictures of the subjects]
@@
    What affected your choice? [Text Answer]
@@
    Your familiarity with photography
    How often do you take photos of yourself or others in your daily life? [Daily or Weekly, Monthly, Every few Months, Yearly or Never]
  8. 7.
    Do you use any software to edit your photos? [Yes/No]
  9. 8.
    Do you spot AI (artificial intelligence) generated pictures in your daily life? [Yes/No]
  10. 9.
    Have you ever used AI to create pictures? [Yes/No]

### v2

@@ 4.2.2 Landmarks Image
For practical use, users often struggle to understand how face positioning in the landmark image maps to the output. This disconnect often leads to dissatisfaction, even though the issue stems from suboptimal conditioning input.
Figure 6: Two-Step Generation Pipeline. Initial outputs use a keypoints image ($s_{kpts}$) from another identity, often reducing facial similarity. Replacing $s_{kpts}$ with a prior output of the subject improves identity preservation while retaining pose. Using four reference images offers a good trade-off, as demonstrated in Appendix 11. Despite the ease-of-use in downstream applications, this limitation motivates our face replacement method (Section 4.2.2) for greater control.
To address this limitation, we propose two solutions: 2-step generation and face replacement.
In 2-step generation (Figure 6), we collect subject reference images $(s_{1},\dots,s_{n})$ and a separate image representing the desired pose and composition $s_{kpts}$. These are used as reference images and the keypoints image, respectively. While the resulting output $s_{out}$ is generally satisfactory, using facial landmarks from one person to generate another reduces facial similarity due to structural differences in the five keypoints (eyes, nose, mouth). We hypothesize this stems from imbalanced conditioning weights. Performance improves when replacing $s_{kpts}$ with a previously generated image of the subject, yielding better facial similarity while maintaining compositional control.
In face replacement, users interact with a simple tool to manipulate (move/rotate/resize) their cropped face on a canvas matching the diffusion model’s output dimensions. This approach eliminates the similarity issues caused by using another person’s facial landmarks. However, the method performs poorly when none of the reference images show the subject facing the camera (deviations $>30\degree$). User satisfaction was higher with this approach compared to 2-step generation, which we attribute to increased interactivity and faster generation times.
@@ 5 Survey
We conducted the survey to evaluate the viability of AI-generated portraits for professional use and to compare the performance of DreamBooth and InstantID in generating realistic headshots. A pilot study was conducted prior to the main survey to ensure the clarity and relevance of the questions. Numerical data on answers can be found in Suppl. 13 and questionary can be found in Suppl. 14.
A total of 97 participants, comprising current white-collar workers as well as students pursuing careers in white-collar fields, took part in the online survey. The participant pool covered ages 18 to 74, with $12.4\%$ of respondents in the younger 18–24 bracket, but only $13.4\%$ in the 25–44 age range. Overall, the sample skewed older, with $74.2\%$ aged 45 or above. Gender was nearly evenly split between male and female, with no representation of diverse identities and several non-disclosures. Among the respondents, 81 self-identified as White, 3 as Asian, 1 as Hispanic, and 12 chose not to disclose their ethnicity. As the sample was predominantly White and older, findings may not fully generalize to younger or more diverse populations.
Overall Performance of Generated Portraits Portraits generated by DreamBooth and InstantID performed similarly across multiple aspects, including overall quality, facial detail clarity, identity preservation, perceived level of editing, and background quality. Using high-quality subject datasets led to slightly better results in most categories, except for “Editing,” where participants indicated familiarity and acceptance of traditional Photoshop-enhanced portraits.
Method Preferences A slightly higher percentage of participants ($\approx$4%) preferred the standardized portraits from InstantID over the more flexible outputs of DreamBooth. InstantID was often perceived as more professional, likely due to its consistent “Photoshopped look,” which resonated with a broader audience. Open-ended responses highlighted diverse preferences, with participants emphasizing factors such as lighting, pose, angle, expression, detail, color, and background.
Facial Similarity DreamBooth demonstrated superior facial similarity between real images individuals and their generated portraits compared to InstantID. More participants identified InstantID images as depicting a different person than the reference. DreamBooth consistently maintained a higher level of facial similarity across both high- and low-quality subject datasets.
Noticing AI Generations Most white-collar workers struggled to identify AI-generated headshots when not explicitly prompted, often focusing on well-known but absent flaws commonly associated with AI generation. Among a subset of participants $(n=77)$ who regularly notice AI-generated images in daily life, the generated portraits blended well with conventional studio photographs. However, participants who actively use AI for image creation $(n=29)$ demonstrated better identification skills. This group was more likely to recognize DreamBooth images as AI-generated, possibly due to DreamBooth’s popularity, while InstantID generations, being more niche, had a near $/50$ 50 chance of being identified as AI.
@@ 6 Discussion
Our experiments offer insights into augmentation strategies for improving facial resemblance in personalized text-to-image generation using DreamBooth and InstantID. While classical augmentations are common in deep learning, applying them to few-shot personalization can yield undesirable results. Geometric transformations like flipping and rotation disrupted learning due to face asymmetry and artifacts. Color jittering caused erratic generations by associating color shifts with DreamBooth’s rare token. Background augmentations with U2-Net introduced segmentation imperfections, especially around hair, which SDXL learned. Replacing backgrounds with patterns or studio backdrops also degraded image quality. Auto color adjustment with Adobe Lightroom improved color grading.
@@
FaceDistance provided a quantitative measure of facial similarity but became less useful for hyperparameter tuning once a certain fidelity level was reached. A user survey among white-collar workers showed that both DreamBooth and InstantID performed similarly in quality, clarity, identity preservation, editing, and background. A slight preference emerged for the “Photoshopped look” of InstantID portraits. While DreamBooth achieved better facial similarity, many participants struggled to distinguish AI-generated images from real ones, particularly those unfamiliar with AI tools. Users actively engaged in AI image creation were more likely to identify DreamBooth images as synthetic, possibly due to its higher popularity.
InstantID’s effectiveness depends on reference image quality and diversity. Using multiple references (around four) improved similarity by enriching information for the Face Encoder. Facial landmarks strongly influenced pose and composition, sometimes overriding text prompts. We explored 2-step generation and interactive face replacement to enhance control, with the latter showing higher user satisfaction. Rotational and shape-altering augmentations were ineffective, and background modifications reduced similarity. Traditional upscaling worked well for low-resolution images, whereas neural network-based upscaling introduced artifacts.
@@ 7 Limitations
We acknowledge that the survey sample is predominantly White and includes mostly older age brackets, which may introduce bias and limit the generalizability of the findings to more ethnically diverse and younger populations.
FaceDistance metric exhibits limited sensitivity, primarily supporting coarse discrimination between low- and high-quality generations rather than capturing nuanced, fine-grained differences. FaceDistance becomes less informative for hyperparameter tuning once generation quality surpasses a baseline level of fidelity. Furthermore, it does not effectively evaluate recontextualization or provide a holistic measure of personalization performance, limiting its utility in tasks where those dimensions are critical.
8 Future Work
Future work should improve FaceDistance’s sensitivity for subtle distinctions and expand its scope beyond facial similarity to include recontextualization and overall image quality. Integrating human preferences will better align evaluations with subjective aesthetics.
Given our goal of enhancing facial fidelity, we plan to adopt LPIPS [28] in future evaluations, as it better captures perceptual differences relevant to visual realism. In contrast, we find DINO [7] is less suitable, emphasizing semantic similarity over identity or fine-grained fidelity.
Future work should consider ethical risks such as misuse, consent, identity theft, and the broader societal impact of personalized generation technologies.
@@ 9 Conclusion
For practical implementation, we establish that using approximately four reference images optimizes InstantID performance, and techniques like 2-step generation and interactive face replacement enhance pose control and user satisfaction. These findings support the viability of generative augmentation as a practical approach for reducing dependency on real data in personalization tasks.
@@
10 FaceDistance in Practice
In downstream applications such as AI-generated portraits, it is common practice to generate large batches of images ($200+$) to ensure sufficient variety and quality. However, this brute-force approach places a significant burden on users, who must manually review and select from hundreds of samples, leading to inefficiency and cognitive overload.
By leveraging FaceDistance to filter and rank images based on their similarity to the reference face, we can greatly reduce this friction. This enables early stopping once enough high-quality matches are found or allows prioritization of the top-ranked images, streamlining the selection process and enhancing overall user satisfaction.
FaceDistance rankings exhibit a strong correlation with those from commercial platforms (Figure 7), demonstrating that our metric effectively captures perceived facial similarity. This suggests FaceDistance captures a fundamental aspect of facial similarity that underlies the ranking criteria used by these platforms, highlighting its novelty and potential as a core component in automated facial assessment.
Figure 21 further shows how samples grouped by FaceDistance correspond to visual likeness, demonstrating FaceDistance as a reliable and interpretable metric for selection that enhances efficiency and satisfaction.
Figure 7: Correlation between FaceDistance and ranking from a commercial website for 229 AI-generated portrait samples. Due to proprietary considerations, we do not disclose the identity of the commercial website.
@@ 11 FaceDistance of InstantID generations
Our experiments with InstantID show that using a single reference image leads to high variance in output quality and inconsistent facial similarity (Figure 8). In contrast, multiple reference images significantly improve facial consistency (Figure 9). Analysis of the Vacation-Anna dataset reveals that while the improvement from 1 to 2 images is substantial, the differences between 2, 4, and 8 images are subtle and often imperceptible without direct side-by-side comparison. The FaceDistance metric supports this, with the distribution shifting leftward as more reference images are used, indicating greater overall similarity.
Figure 8: Distribution of facial similarity metrics for 8 different input subjects using single reference images, with the y-axis sorted by FaceDistance (meaning Img_3 has the lowest FaceDistance within, Img_5 has the highest).
@@
Figure 10: Illustrative Images of Figure 8. Generated using the 2-step method (Section 4.2.2).
12 Motivating Alternatives to InstantID
@@ 13 Comparative Survey Analysis of Portrait Generation Methods
Our analysis reveals that generated portraits perform similarly across individual aspects, likely due to the SD model employed (Figure 13). Good subject datasets for Dreambooth yield more natural results, though user preferences trend toward Photoshop-edited appearances. Further investigation shows that “good” datasets correlate with higher employee agreement on preferred methods, with approximately 4% more participants favoring standardized InstantID portraits for their professional, photoshopped aesthetic (Figure 14). When assessing facial similarity between real and generated images, Dreambooth consistently outperformed InstantID regardless of dataset quality, with fewer participants identifying Dreambooth generations as different individuals (Figure 15).
White-collar workers demonstrated limited ability to distinguish AI-generated images, often searching for well-known AI flaws absent from our high-quality generations. Among the subset of participants ($n=77$) who reported noticing AI-generated images in daily life, our generations integrated seamlessly with non-generated studio images, with only the “Real (Daniel)” image—featuring strong contrast and sharp features—dividing opinion nearly equally (Figure 17, Figure 16).
@@
Figure 14: User Study Results: Participant Choices for ’Is Professional?’ and ’Preferred Method’ per AI Generation Technique.
Figure 15: User Study Results: Participant Similarity Ratings (’Is the Same Person?’) for Generated Images by Method
Figure 16: User Study Results: Participant Responses to ’Is This Picture AI Generated?’ per Image $(n=77)$
Figure 17: Sample images from Real/Fake Questions. (Results in Figure 16)
14 Survey Questionnaire
The following is the complete questionnaire used in our user study. Answer options are shown in italic, while section titles and key instructions are marked in bold. Where applicable, questions were paired with image descriptions to provide visual context.
@@ 1. Title: Survey on Studio Portraits
    Brief introductory text shown to participants.
@@
    Rate the skill level of the photographer
    ((This section was repeated for each of four subjects. Each subject had four generated portraits shown in a grid (e.g., Figure 11, Figure 12) along with a brief description of their professional background (e.g., “marketer”, “researcher”, “nurse”). Use of AI is not disclosed to participants, yet.))
@@
      How would you rate the overall quality?
      [Really Bad] 1–5 [Really Good]
@@
      Are the facial details clear and well-defined?
      [No] 1–5 [Yes]
@@
      [Completely Different] 1–5 [Exactly the Same]
@@
      [No Editing] 1–5 [Heavily Edited]
@@
      Rate the quality of the background in the headshot.
      [Poor] 1–5 [Excellent]
@@
      Would you expect to see these photos in a professional context (e.g., LinkedIn)? [Yes/No]
@@
      Is there anything you dislike about these pictures?
      [Text Answer]
@@
    Photographer Preference
    If you liked at least one image, select the square near it.
    [Four image grids from above]
@@
    Similarity of Real-Life Pictures and Portraits
    Compare how similar the person in the everyday photo (left) is to the person in the portrait (right).
    How would you rate the similarity of the person?
    [Not the Same] 1–3 [The Same]
@@
    Detection of AI Use
    Some images may have been created using AI. Do not revise earlier responses.
    Please select any photos you believe were AI-generated.
    [Four subject images]
@@
    What influenced your choice? [Text Answer]
@@
    Participant Background: Photography Experience
    1. (a)
      How often do you take photos of yourself or others?
      [Daily or Weekly, Monthly, Every Few Months, Yearly or Never]
    2. (b)
      Do you use any software to edit your photos?
      [Yes/No]
    3. (c)
      Do you frequently encounter AI-generated images in daily life? [Yes/No]
    4. (d)
      Have you ever used AI tools to create images?
      [Yes/No]
We thank all participants for their time and contributions.
15 Flickr-Suits-XL and Flickr-Portraits-XL Datasets
We have collected a new dataset of people in suits, Flickr-Suits-XL (FSXL), consisting of 1208 high-quality images (Figure 18). A noticeable gender imbalance is present in the FSXL dataset, with male subjects being the majority.
Figure 18: Samples from the Flickr-Suits-XL (FSXL) dataset. Images are center-cropped in this preview.
These images were obtained by crawling Flickr for 5687 images (thus inheriting all the biases of that website) under permissive licenses. We kept images with 1 visible face with MTCNN [27], aligned [12] and cropped the images (to the closest aspect ratio in SDXL training resolutions [17] i.e. $\approx 1$MP), maintaining the detected face centered horizontally and $\frac{1}{3}$ from top vertically.
We have made the dataset publicly available at
https://github.com/KorayUlusan/fsxl-dataset
Flickr-Portraits-XL (FPXL) follows the same procedure as FSXL but uses “in-the-wild” raw images of FFHQ dataset [11], inheriting all the biases. (Figure 19) We thank the authors of the FFHQ dataset.
We have made the dataset publicly available at
https://github.com/KorayUlusan/fpxl-dataset
Figure 19: Samples from the Flickr-Portraits-XL (FPXL) dataset. Images are center-cropped in this preview (making it look like FFHQ).
One of the motivations in writing this paper was to create LinkedIn style portraits. Collecting such images was a natural step forward.
16 Investigating Alternative Regularization for DreamBooth: A Case Study on Reconfirming Class Prior Preservation Principles
DreamBooth [20] effectively personalizes large text-to-image diffusion models, enabling the synthesis of novel renditions of specific subjects from just a few input images. A critical element for its success is the autogenous class-specific prior preservation loss (PPL). This loss is specifically designed to counteract two common issues during fine-tuning: “language drift,” where the model forgets how to generate subjects of the target class, and the risk of reduced output diversity. DreamBooth mitigates these by supervising the model with its own generated samples, thereby leveraging and retaining the semantic prior embedded within the model itself. This approach is crucial for preventing overfitting and ensuring the personalized subject can be generated in diverse contexts, poses, and views while maintaining its identity and the broader class knowledge.
In our study, complementing our primary research on augmentations, we investigated an alternative regularization strategy: utilizing external image collections instead of the model’s self-generated samples for prior preservation. We used Flickr-Portraits-XL (FPXL) and and Flickr-Suits-XL (FSXL) datasets (Appendix 15) to asses whether they can serve as regularization images to preserve the class prior during DreamBooth fine-tuning, akin to the objective of DreamBooth’s PPL.
However, our experiments revealed a significant and unexpected outcome: using FPXL and FSXL as regularization datasets fundamentally destroyed the class word prior. After only a few training steps, the DreamBooth model could no longer properly generate images of “a [class word]” (e.g., “a man” or “a woman”). This failure occurred because the distributions of these external datasets diverged significantly from the Stable Diffusion model’s inherent prior for the class word. Additionally, differences in camera field of view and cropping in FPXL and FSXL images—compared to the self-generated images typically used for regularization—exacerbated this divergence, especially with a fast-converging learning rate.
The consequences of this prior destruction were severe: training either diverged, resulting in heavily deformed objects, or the model severely overfitted the subject dataset. Figure 20 illustrates the model’s behavior prior to the breakdown of the class prior. A distinct qualitative indicator of a destroyed class prior was the appearance of a purple tint in generated class-word images (e.g., “a $[V]$ man”).
Figure 20: Artifacts in DreamBooth training by using external image collections as regularization datasets.
This negative result provides strong empirical support for DreamBooth’s autogenous prior preservation loss [20]. Our findings highlight that drastically altering the class prior using external regularization datasets is generally infeasible. Instead, finetuning the model remains the preferred method to shift the class prior. This case study thus serves as a qualitative ablation, illustrating the pitfalls of alternative regularization strategies and reaffirming DreamBooth’s PPL [20].
Real Min (Best) Q1 (0.182) Q2 (0.200) Q3 (0.229) Max (Worst)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 21: Visual Comparison of Real and Generated Faces Across Similarity Bins. Columns represent increasing face distance (left to right), while rows display varied samples. Faces are sorted by similarity using FaceDistance, ranging from 0.109 to 0.312.